3.2 4.7 3.8 4.3 3.5 5.0 4.3 5.2 5.1 4.2





1.Constuct a frequency destribution mode.





2.Determined the interval where k=10.





3.Compute for mean,median ,mode ,variance and standard deviation of grouped date

1
Expert's answer
2021-12-30T17:34:07-0500

3.2 4.7 3.8 4.3 3.5 5.0 4.3 5.2 5.1 4.2 1.8 5.1 5.8 5.0 5.6 4.5 4.8 4.0 2.7 3.8 4.2 4.1 2.7 5.6 6.0 3.9 4.2 3.4 5.5 3.8 5.1 4.6 5.8 4.3 2.5 5.8 4.5 4.9 4.7 4.4 2.4 3.4 4.9 4.3 3.7 4.9 2.7 3.0 4.2 4.0 3.6 3.9 3.1 3.9 5.3 4.4 3.7 3.1 2.8 3.8


Least to Greatest Value:

1.8, 2.4, 2.5, 2.7, 2.7, 2.7, 2.8, 3, 3.1, 3.1, 3.2, 3.4, 3.4, 3.5, 3.6, 3.7, 3.7, 3.8, 3.8, 3.8, 3.8, 3.9, 3.9, 3.9, 4, 4, 4.1, 4.2, 4.2, 4.2, 4.2, 4.3, 4.3, 4.3, 4.3, 4.4, 4.4, 4.5, 4.5, 4.6, 4.7, 4.7, 4.8, 4.9, 4.9, 4.9, 5, 5, 5.1, 5.1, 5.1, 5.2, 5.3, 5.5, 5.6, 5.6, 5.8, 5.8, 5.8, 6


1.




2.

interval where cumulative frequency k=10:

class 3 - 4


3.

mean:

"\mu=\sum x_if_i=4.22"

where xi is midpoint of class,

fi is frequency


for Median Class:

value of (n/2)th observation = value of (60/2)th observation =

= value of 30th observation

The median class is 4-5

median:

"m=L+\frac{n/2-cf}{f}c=4+\frac{30-24}{22}=4.27"

where

L=lower boundary point of median class =4

n=Total frequency =60

cf=Cumulative frequency of the class preceding the median class =24

f=Frequency of the median class =22

c=class length of median class =1


for Mode Class:

maximum frequency is 22.

The mode class is 4-5

mode:

"M=L+\frac{f_1-f_0}{2f_1-f_0-f_2}c=4+\frac{22-17}{2\cdot22-17-13}=4.36"


Variance:

"\sigma^2=\frac{\sum x^2_if_i-\sum x_if_i/n}{n}=1.04"


standard deviation:

"\sigma=\sqrt{\frac{\sum x^2_if_i-\sum x_if_i/n}{n}}=1.018"